OpenStudy (anonymous):

Which equation shows the substitution method being used to solve the system of linear equations? x + y = 6 x = y + 2 A)) x + y = y + 2 B)) x + (y + 2) = 6 C)) x = (x – 6) + 2 D)) (y + 2) + y = 6

TheSmartOne (thesmartone):

So basically they want to know what is the next step in solving that substitution. So we have x=y+2 we want to substitute the value we get for x into x+y=6

TheSmartOne (thesmartone):

@CrazyCountryGirl

OpenStudy (anonymous):

its c, becuase you can substute the value of x into the equation

TheSmartOne (thesmartone):

Actually C is wrong.

OpenStudy (anonymous):

explain @TheSmartOne

TheSmartOne (thesmartone):

x + y = 6 x = y + 2 since x= y+2 we could plug that into x+y=6 So x+y=6 (y+2)+y=6

OpenStudy (anonymous):

c. x= (x-6)+2

OpenStudy (anonymous):

your solution of course will work @TheSmartOne

OpenStudy (anonymous):

mine won t work!

TheSmartOne (thesmartone):

What I just proved -_-

OpenStudy (anonymous):

because I have x =x

OpenStudy (anonymous):

and that is why you r @TheSmartOne